Circular No. 7004

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SGR 1900+14 K. Hurley, Space Sciences Laboratory, University of California, Berke-

ley; T. Cline, Goddard Space Flight Center, NASA; and E. Mazets and S. Golenetskii, Ioffe Institute, report: “The extraordinary KONUS burst reported by Cline et al. (IAUC 7002) has been observed and triangulated with Ulysses. The annulus of possible arrival directions is indeed consistent with the position of this soft repeater (IAUC 7001). The 5.1-s periodicity is clearly detected in the Ulysses data for 300 s following the burst and in another event that followed it ∼ 6 hr later. A light curve may be viewed

at http://ssl.berkeley.edu/ipn3/newdata.html.”

NOVA IN NGC 221 E. Halderson, T. Shefler, M. Modjaz, J. Y. King, W. D. Li, R. R.

Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent nova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The nova was found and confirmed on unfiltered images taken on Aug. 31.4 (mag ∼ 16.6) and Sept. 1.3 UT (mag ∼ 16.5). The object is located at α = 0^{h}42^{m}39^{s}.29, δ = +40^{o}51′10′′.3 (equinox 2000.0), which is ∼ 28′′.5 west and 44′′.7 south of the nucleus of NGC 221 (the satellite galaxy of M31). A KAIT image obtained on Aug. 27.3 UT did not show a star at the position of the nova (limiting mag ∼ 19.0).

S. Benetti, Telescopio Nazionale Galileo, La Palma; and T. Pursimo and P. Heinamaki, Tuorla Observatory, confirm the presence of a new object of magnitude R = 15.87 ± 0.03 located 28′′.5 west and 45′′.1 south of the nucleus of NGC 221, on an R frame taken with the Nordic Optical Telescope at La Palma on Sept. 1.12 UT.

COMET C/1998 M5 (LINEAR)

Ephemeris extension to IAUC 6973 (elements MPC 32169):

1998TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Sept. 4 21^{h}02.^{m}28 +39^{o}28.^{′}6 1.758 2.497 126.^{o}9 18.^{o}9 11.2
9 20 45.44 +40 04.3 1.748 2.455 123.8 19.9 11.1
14 20 28.93 +40 23.3 1.747 2.414 120.1 21.1 11.0
19 20 13.19 +40 27.0 1.753 2.373 116.1 22.3 11.0
24 19 58.56 +40 17.7 1.765 2.333 111.9 23.5 10.9
1998 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green